# Zero-Knowledge Federated Learning: A New Trustworthy and Privacy-Preserving Distributed Learning Paradigm

### v2

@@ Abstract
Index Terms:
Federated Learning, Zero-Knowledge Proof, Trust and Security, Privacy-Preserving.
@@ I Introduction
Zero-Knowledge Proof (ZKP) [3] is a cryptographic technique that can be integrated into FL to bolster trust among participants—a paradigm known as Zero-Knowledge Federated Learning (ZK-FL). However, research on ZK-FL is still in its early stages, and there is a lack of a systematic classification and comprehensive summary of the related work. In this paper, we introduce a structured ZK-FL framework that categorizes the technical roles of ZKP within each module of FL, and we classify, analyze, and summarize existing ZK-FL studies accordingly. At the same time, we propose a novel FL algorithm that applies ZKP during the client selection process, named the Verifiable Client Selection FL (Veri-CS-FL) algorithm. In this approach, while local models are trained across various participating clients, cosine similarity metrics between these models and a benchmark model are used for evaluation. The benchmark model is derived by training on the central server using its proprietary data. During metric generation, ZKP is employed to produce proofs that verify the accuracy of the metric calculation. The server then only selects and aggregates the local models with higher-ranking, validated metrics. Lower-ranking, poorly performing local models no longer need to be uploaded to the server. By incorporating ZKP, our approach not only enhances the evaluation and communication efficiencies of local models through transparent evidence of their quality but also fosters increased trust among clients.
@@ II-A Federated Learning
Existing FL algorithms often operate under the assumption that both clients and the server are inherently trustworthy, an assumption that may not hold in practical distributed scenarios. Addressing the challenges of safeguarding data privacy and mitigating the impact of Byzantine nodes during the FL process is therefore critical. This article focuses on the zero-knowledge proof based strategies proposed to prevent malicious behaviors from both servers and clients.
@@ II-B Zero-Knowledge Proof
ZKP is a cryptographic technique that proves a statement’s validity without revealing any private information about the statement. There are several types of ZKP algorithms [4]. Among them, zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) is considered to be the most practical. A zk-SNARK algorithm is usually represented by an arithmetic circuit that consists of the basic arithmetic operations of addition, subtraction, multiplication, and division. The three algorithmic components of a zk-SNARK algorithm are given:
@@
    $(PK,VK)\leftarrow KEYGEN(1^{\lambda},R)$ is the key generation algorithm that generates the proving key $PK$ and the verification key $VK$ by using a predefined security parameter $\lambda$ and an arithmetic circuit representation $R$.
@@
    $\pi\leftarrow PROVE(PK,x,W)$ is the proof generation algorithm that generates a proof $\pi$ based on the proving key $PK$, the input $x$ and the witness $W$.
@@
    $1/0\leftarrow VERIFY(VK,x,\pi)$ is the proof verification algorithm that outputs a decision to accept or reject $\pi$ using $VK$, $x$ and $\pi$ as the input.
The proving key $PK$ and the verification key $VK$ generated by the $KEYGEN$ algorithm are treated as the public parameters pregenerated by an authority. The $PROVE$ algorithm is executed by the prover, and the $VERIFY$ algorithm is executed by the verifier. Witness $W$ is the secret owned by the prover that he/she does not want to reveal to others and yet wants to prove that he/she knows the secret. The technical advantages of zk-SNARK inlcudes its short proof size, fast proof verification, non-interactive proving [4]. In this article, we consider use zk-SNARK to generate zero-knowledge proofs to provide trusts for our Veri-CS-FL algorithm.
@@ III-A1 Client Identity Trust
  • ZKP Solution: ZKP can verify client identities without exposing sensitive information. For instance, [5] introduces a lightweight authentication framework for blockchain-based FL, using ZKP to verify client identities. Similarly, [6] applies ZKP to protect client identity privacy in vehicular edge computing.
@@ III-A2 Local Data Trust
  • ZKP Solution: ZKP can ensure the authenticity and integrity of local training data by requiring clients to create cryptographic commitments to their data and generate proofs that the committed data was used for training. This approach preserves data confidentiality while ensuring reliability.
@@ III-A3 Local Training Trust
  • ZKP Solution: ZKP can verify that local training processes adhere to predefined protocols. Works such as [7, 8, 9, 10, 11] utilize ZKP to ensure the integrity of local training computations, preventing low-quality or malicious models from contaminating the global model.
@@ III-B2 Model Aggregation Trust
  • ZKP Solution: ZKP can ensure trustworthy aggregation by generating verifiable proofs that the server adheres to agreed-upon protocols. Works such as [7, 12] demonstrate how ZKP can validate the correctness and fairness of the aggregation process, ensuring that client contributions are fairly incorporated into the global model.
@@
Stage Trust Issue Solution Reference
Local Model Training Stage Client Identity Trust Ensuring the Authenticity of Client Identities [5], [6]
Local Data Trust Ensuring the Integrity and Accuracy of the Training Data  —
Local Training Trust Ensuring the Integrity and Accuracy of the Model Updates [7], [8], [9], [10], [11]
Global Model Aggregation Stage Model Evaluation Trust Ensuring Integrity of Performance Metric Evaluations [11] and this work
Model Aggregation Trust Ensuring Trustworthy Model Aggregation [7], [12]
TABLE I: Summary of the existing ZK-FL solutions
@@ IV-A Threat and Trust Model
  • Malicious clients may launch poisoning attacks by submitting falsified performance metrics of their local models, with the intent of disrupting the overall learning process.
Under this threat and trust model, updating the global model with poorly performing local models uploaded from malicious clients could severely compromise FL system performance. To mitigate this risk, the server must access each uploaded model’s performance and select those local models that meet predefined performance standards to aggregate. This process is referred to as client selection. While a straightforward method for client selection is to have clients report their model performance metrics and then to select, this self-reporting is unreliable due to the inherent mistrust between the server and clients, as well as the risk of malicious behavior. Consequently, the server should independently evaluate each client’s local model before selection. However, this approach introduces two significant drawbacks:
@@ IV-B Veri-CS-FL
In this section, we propose our ZK-FL algorithm, i.e., the Veri-CS-FL algorithm, to achieve the trusted model evaluation at the client-side and client selection at the server-side without re-evaluating the local models submitted by the clients for verifications. Veri-CS-FL is also based on the FedAvg algorithm framework, and it leverages local models’ performance metrics as the criterion to select the clients’ local models with satisfied performances for aggregation. Other than re-running all the clients’ local models to compute their performance metrics at the server side, Veri-CS-FL integrates ZKP to validate the accuracy of the computed performance metrics in a communication-and-computation-efficient and privacy-preserving way. The specific processes of Veri-CS-FL are described as follows.
On the server side, Veri-CS-FL designates a small and clean dataset $D_{s}$ as the root dataset, enabling the server $S$ to train a benchmark model to evaluate the performances of the clients’ local models. At the start of the FL training process, the key generation algorithm of zk-SNARK is executed at a trusted third party to generate the proving key ($PK$) for all clients, who act as provers, and the verification key ($VK$) for the server, who acts as the verifier: $(PK,VK)\leftarrow KEYGEN(1^{\lambda},R)$, where $R$ denotes the circuit that models the computation process for evaluating model performance. At the start of the $t$-th FL training round, all entities, including the server and the clients, access the global model, $w^{t-1}$, trained in the previous FL training round. Then, the following training steps are conducted between the server and the clients in the $t$-th FL training round:
@@
    Each client $C_{k}$ trains an updated local model $w_{k}^{t}$ from the last round global model $w^{t-1}$ utilizing its local dataset $D_{k}$, $k\in{1,2,\cdots,K}$ [^1] We assume that the local model training stage is conducted honestly by all clients, which is further guaranteed through a ZKP algorithm, as demonstrated in [7, 8, 9, 10, 11]. We only consider the model evaluation trust in the design of Veri-CS-FL. Subsequently, each client computes a performance metric of the updated local model $w_{k}^{t}$ by comparing the updated local model $w_{k}^{t}$ with the benchmark model $w_{s}^{t}$, and employs the proof generation algorithm of zk-SNRAK to generate a zero-knowledge proof $\pi_{k}$ for the performance metric of $w_{k}^{t}$: $\pi_{k}\leftarrow PROVE(PK,x,W)$, where the public input $x$ is the benchmark model of the server $w_{s}^{t}$ together with the hash of the updated local model $h_{k}^{t}=\text{hash}(w_{k}^{t})$: $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$, and the witness $W$ is the updated local model $w_{k}^{t}$: $W=w_{k}^{t}$.
@@
    Each client $C_{k}$ transmits the proof $\pi_{k}$ for the updated local model’s performance metric, the hash of their updated local model $h_{k}^{t}$, and the corresponding performance metric to the server $S$, other than transmitting the updated local model $w_{k}^{t}$ to the server at this stage.
@@
    The server $S$ collects the performance metrics of the clients’ local models and their corresponding proofs, $\pi_{k}$, $k\in{1,2,\cdots,K}$. The server first verifies all the proofs using the proof verification algorithm of zk-SNARK: $1/0\leftarrow VERIFY(VK,x,\pi)$ with $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$ and $\pi=\pi_{k}$, and proceeds to rank the corresponding performance metrics whose proofs have been successfully verified. The performance metrics whose proofs cannot pass the verification are not used and ranked. The server $S$ then selects the top $N$ clients based on their ranked performance metrics into the candidate set $\mathcal{C}$ and notifies the respective clients in this set to upload their updated local models.
@@
    The server $S$ collects the updated local models sent from the $N$ clients in the candidate set $\mathcal{C}$. The server first computes the hash of the updated local model sent by each selected client in step 5: $\widetilde{h}_{k}^{t}=\text{hash}(w_{k}^{t})$, and compares the computed hash with the hash, $h_{k}^{t}$, sent by the client in step 3. If the two hashes match, the server $S$ will labels this client as a valid candidate; if the two hashes differs, the server $S$ will remove this client $C_{k}$ from the candidate set $\mathcal{C}$. After this hash verification operation for each client $C_{k}\in\mathcal{C}$, the server $S$ aggregates these updated local models of the clients still in the candidate set $\mathcal{C}$ according to the predetermined rule to derive the updated global model $w^{t}$ for the $t$-th round and then distribute $w^{t}$ to all the clients.
The above processes of Veri-CS-FL iterate continuously until the final global model converges or the predetermined number of epochs reaches. Some remarks about Veri-CS-FL are given below.
0:  Each client $C_{k}$ possesses the proving key $PK$ of zk-SNARK and the local dataset $D_{k}$ , $\forall k\in\{1,2,\cdots,K\}$ while the server $S$ holds the verifying key $VK$ of zk-SNARK and the small and clean dataset $D_{s}$.
@@
9:        generates $\pi_{k}^{t}$ of $w_{k}^{t}$: $\pi_{k}^{t}\leftarrow PROVE(PK,x,W)$ with $h_{k}^{t}=\text{hash}(w_{k}^{t})$, $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$ and $W=w_{k}^{t}$;
[…]
@@
14:         verifies all the proofs: $1/0\leftarrow VERIFY(VK,x,\pi)$ with $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$ and $\pi=\pi_{k}$;
@@
22:         compares $\widetilde{h}_{k}^{t}$ with $h_{k}^{t}$ (line 9): if $\widetilde{h}_{k}^{t}\neq h_{k}^{t}$, reomves the corresponding $C_{k}$ from $\mathcal{C}$, $\forall k\in\{1,2,\cdots,N\}$.
@@
    In Veri-CS-FL, each client needs to submit the performance metric of their local model to the server. The server then uses the performance metrics to select local models with satisfied performances, and uses the training sample sizes as the weights in the weighted aggregation method of global models. There are many different performance metrics that can be used to characterize the performance of a local model. For example, the loss function of the model on the test dataset, the KL divergence of the output predicted by the model on the test dataset. In Veri-CS-FL, we propose to adopt the cosine similarity between the benchmark model and the local model [13] as the performance metric:
    $\text{cos}_{k}^{t}\overset{\triangle}{=}{\text{cos}}\left(w_{k}^{t},w_{s}^{t}\right)=\frac{w_{k}^{t}\cdot w_{s}^{t}}{|w_{k}^{t}||w_{s}^{t}|}$ (1)
    When utilizing the benchmark model as a reference, the larger the cosine similarity between the local model and the benchmark model, the better the performance of the local model. The computation of cosine similarity can also be easily translated to the arithmetic circuit of zk-SNARK, enabling the computation of the zero-knowledge proof for it.
@@
    By generating the zero-knowledge proofs for the performance metrics of local models at the client-sides, and validating the proofs at the server side, Veri-CS-FL can achieve the following advantages: i) Communication-efficient: the local models with poor performances will not be communicated between the clients and the server. This can save much uplink communication bandwidth, particularly when the models are large. ii) Computation-efficient: the server has no need to re-run all the local models to evaluate their performances. The verification of zero-knowledge proofs in zk-SNARK is very efficient, e.g., the zk-SNARK algorithm of Groth16 [14] has a fixed verification complexity and fixed proof size regardless of the sizes of public inputs and private witness. iii) Trusted collaboration: The client transmits the performance metrics and its zero-knowledge proofs to the server, and if the server feels that its updated local model has satisfactory performance, it asks the client for the updated local model. As a result, the client and server can agree on the model training contribution. It avoids that after the client passes the updated local model to the server in advance, the server refuses to admit that the client provided a local model with satisfactory performance in this round of training, and maliciously devalues the actual contribution of the client.

### v3

@@ I Introduction
To address these challenges, zero-knowledge proof (ZKP) [3] based federated learning (ZK-FL) has emerged as a powerful solution, particularly in cross-organizational or other high-stakes settings (e.g., finance, healthcare, and autonomous driving) where mutual trust between the server and clients is limited and strong auditability and computational integrity guarantees are required. ZK approaches specifically address threats related to client malfeasance that other privacy methods (like differential privacy or homomorphic encryption) often miss, such as free-riding attacks and model poisoning via quality verification. ZK-FL is particularly suitable for scenarios requiring verifiable correctness, hardware independence (unlike Trusted Execution Environments which rely on specific hardware), and communication efficiency.
The contribution of this paper is twofold, as summarized below:
  • $\bullet$
    While ZKP can be integrated into FL to bolster trust among participants, research on ZK-FL is still in its early stages, and there is a lack of a systematic classification and comprehensive summary of the related work. In this paper, we thus first introduce a structured ZK-FL framework that categorizes the technical roles of ZKP within each stages of FL (e.g., local training, model evaluation, aggregation), and we classify, analyze, and summarize existing ZK-FL studies accordingly.
  • $\bullet$
    Based on the structured ZK-FL framework, we then propose a novel FL algorithm that applies ZKP during the client selection process within the “model evaluation” stage, named the Verifiable Client Selection FL (Veri-CS-FL) algorithm. In this approach, while local models are trained across various participating clients, cosine similarity metrics between these models and a benchmark model are used for evaluation. The benchmark model is derived by training on the central server using its proprietary data. During metric generation, ZKP is employed to produce proofs that verify the accuracy of the metric calculation. The server then only selects and aggregates the local models with higher-ranking, validated metrics. Lower-ranking, poorly performing local models no longer need to be uploaded to the server. By incorporating ZKP, our approach not only enhances the evaluation and communication efficiencies of local models through transparent evidence of their quality but also fosters increased trust among clients.
@@ II-A Federated Learning
Existing FL algorithms often operate under the assumption that both clients and the server are inherently trustworthy, an assumption that may not hold in practical distributed scenarios. Addressing the challenges of safeguarding data privacy and mitigating the impact of Byzantine nodes during the FL process is therefore critical. In this context, Byzantine nodes refer to participants that behave arbitrarily or maliciously—ranging from sending incorrect model updates to intentionally disrupting the training protocol—posing a severe threat to system reliability. This article focuses on the zero-knowledge proof based strategies proposed to prevent malicious behaviors from both servers and clients.
@@ II-B Zero-Knowledge Proof
ZKP is a cryptographic technique that proves a statement’s validity without revealing any private information about the statement. Among several types of ZKP algorithms, zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) is considered to be the most practical. A zk-SNARK algorithm is usually represented by an arithmetic circuit that consists of the basic arithmetic operations of addition, subtraction, multiplication, and division. The three algorithmic components of a zk-SNARK algorithm are given:
@@
    $(\mathsf{PK},\mathsf{VK})\leftarrow\mathsf{KEYGEN}(1^{\lambda},R)$ is the key generation algorithm that generates the proving key $\mathsf{PK}$ and the verification key $\mathsf{VK}$ by using a predefined security parameter $\lambda$ and an arithmetic circuit representation $R$.
@@
    $\pi\leftarrow\mathsf{PROVE}(\mathsf{PK},x,W)$ is the proof generation algorithm that generates a proof $\pi$ based on the proving key $\mathsf{PK}$, the input $x$ and the witness $W$.
@@
    $b\leftarrow\mathsf{VERIFY}(\mathsf{VK},x,\pi)$ is the proof verification algorithm that outputs a binary decision to accept ($b=1$) or reject ($b=0$) $\pi$ using $\mathsf{VK}$, $x$ and $\pi$ as the input.
The proving key $\mathsf{PK}$ and the verification key $\mathsf{VK}$ generated by the $\mathsf{KEYGEN}$ algorithm are treated as the public parameters pre-generated by an authority. The $\mathsf{PROVE}$ algorithm is executed by the prover, and the $\mathsf{VERIFY}$ algorithm is executed by the verifier. Witness $W$ is the secret owned by the prover that he/she does not want to reveal to others and yet wants to prove that he/she knows the secret. The technical advantages of zk-SNARK include its short proof size, fast proof verification, and non-interactive proving. In this article, we consider using zk-SNARK to generate zero-knowledge proofs to provide trust for our Veri-CS-FL algorithm. From a privacy perspective, the zero-knowledge property of zk-SNARK ensures that the generated proof $\pi$ reveals nothing about the witness $W$ other than the validity of the statement being proven. This cryptographic guarantee prevents any information leakage regarding the prover’s private data during the verification process.
@@ III-A1 Client Identity Trust
  • ZKP Solution: ZKP can verify client identities without exposing sensitive information. For instance, [6] introduces a lightweight authentication framework for blockchain-based FL, using ZKP to verify client identities.
@@ III-A2 Local Data Trust
  • ZKP Solution: ZKP can ensure the authenticity and integrity of local training data by requiring clients to create cryptographic commitments to their data and generate proofs that the committed data was used for training the model [1]. This approach preserves data confidentiality while ensuring reliability.
TABLE I: Summary of the Existing ZK-FL Solutions.
Stage Trust Issue Solution Reference
Local Model Training Stage Client Identity Trust ①Ensuring the Authenticity of Client Identities [6]
Local Data Trust ②Ensuring the Integrity and Accuracy of the Training Data [1]
Local Training Trust ③Ensuring the Integrity and Validity of Model Updates [5], [13], [8],[2]
Global Model Aggregation Stage Model Evaluation Trust ④Ensuring Integrity of Performance Metric Evaluations [11] and this work
Model Aggregation Trust ⑤Ensuring Trustworthy Model Aggregation [9], [14], [12]
@@ III-A3 Local Training Trust
  • ZKP Solution: ZKP can verify that local training processes adhere to predefined protocols. Works such as [5, 13, 8, 2] leverage ZKPs to verify protocol-compliant local computation, e.g., proving that reported updates are computed from the prescribed model and training procedure (rather than being fabricated or arbitrarily manipulated), while preserving the confidentiality of local data.
@@ III-B2 Model Aggregation Trust
  • ZKP Solution: ZKP can ensure trustworthy aggregation by providing verifiable evidence that the aggregation procedure follows the agreed rule and that malformed contributions are detectable. Works such as [9, 14] demonstrate how ZKPs can be integrated into FL to enable verifiable aggregation under Byzantine behaviors (including decentralized settings), improving transparency and integrity of the aggregation outcome. Additionally, [12] proposes a Byzantine-robust FL framework that embeds ZKPs to ensure verifiable aggregation even under a malicious server, addressing threats in dynamic and non-IID scenarios.
@@ IV-A Threat and Trust Model
  • Malicious clients can engage in poisoning attacks by submitting unreliable models with poor performance, aiming to disrupt the overall learning process.
Under this threat and trust model, updating the global model with poorly performing local models uploaded from malicious clients could severely compromise FL system performance. To mitigate this risk, the server must access each uploaded model’s performance and select those local models that meet predefined performance standards to aggregate. This process is referred to as client selection. While a straightforward method for client selection is to have clients report their model performance metrics and then select, this self-reporting is unreliable due to the inherent mistrust between the server and clients, as well as the risk of malicious behavior. Consequently, the server should independently evaluate each client’s local model before selection. However, this approach introduces two significant drawbacks:
@@ IV-B Veri-CS-FL
In this section, we propose our ZK-FL algorithm, i.e., the Veri-CS-FL algorithm, to achieve the trusted model evaluation at the client-side and client selection at the server-side without re-evaluating the local models submitted by the clients for verifications. Veri-CS-FL is also based on the FedAvg algorithm framework, and it leverages local models’ performance metrics as the criterion to select the clients’ local models with satisfied performances for aggregation. Rather than re-running all the clients’ local models to compute their performance metrics at the server side, Veri-CS-FL integrates ZKP to validate the accuracy of the computed performance metrics in a communication-and-computation-efficient and privacy-preserving way. The specific processes of Veri-CS-FL are described as follows:
On the server side, Veri-CS-FL designates a small and clean dataset $D_{s}$ as the root dataset, enabling the server $S$ to train a benchmark model to evaluate the performances of the clients’ local models. At the start of the FL training process, the key generation algorithm of zk-SNARK is executed at a trusted third party to generate the proving key ($\mathsf{PK}$) for all clients, who act as provers, and the verification key ($\mathsf{VK}$) for the server, who acts as the verifier: $(\mathsf{PK},\mathsf{VK})\leftarrow\mathsf{KEYGEN}(1^{\lambda},R)$, where $R$ denotes the circuit that models the computation process for evaluating model performance. At the start of the $t$-th FL training round, all entities, including the server and the clients, access the global model, $w^{t-1}$, trained in the previous FL training round. Then, the following training steps are conducted between the server and the clients in the $t$-th FL training round:
@@
    Each client $C_{k}$ trains an updated local model $w_{k}^{t}$ from the last round global model $w^{t-1}$ utilizing its local dataset $D_{k}$, $k\in{1,2,\cdots,K}$. [^1] We assume that the local model training stage is conducted honestly by all clients, which is further guaranteed through a ZKP algorithm, as demonstrated in [5, 13, 8, 2]. We only consider the model evaluation trust in the design of VeriCS-FL. Subsequently, each client computes a performance metric of the updated local model $w_{k}^{t}$ by comparing the updated local model $w_{k}^{t}$ with the benchmark model $w_{s}^{t}$, and employs the proof generation algorithm of zk-SNRAK to generate a zero-knowledge proof $\pi_{k}$ for the performance metric of $w_{k}^{t}$: $\pi_{k}\leftarrow\mathsf{PROVE}(\mathsf{PK},x,W)$, where the public input $x$ is the benchmark model of the server $w_{s}^{t}$ together with the hash of the updated local model $h_{k}^{t}=\text{hash}(w_{k}^{t})$: $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$, and the witness $W$ is the updated local model $w_{k}^{t}$: $W=w_{k}^{t}$.
@@
    Each client $C_{k}$ transmits the proof $\pi_{k}$ for the updated local model’s performance metric and the hash of their updated local model $h_{k}^{t}$ to the server $S$. Instead of transmitting the full model immediately, clients first transmit the proof. Only selected clients upload the model subsequently.
@@
    The server $S$ collects the performance metrics of the clients’ local models and their corresponding proofs, $\pi_{k}$, $k\in{1,2,\cdots,K}$. The server first verifies all the proofs using the proof verification algorithm of zk-SNARK: $\mathsf{VERIFY}(\mathsf{VK},x,\pi)$ with $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$ and $\pi=\pi_{k}$, and proceeds to rank the corresponding performance metrics whose proofs have been successfully verified. The performance metrics whose proofs cannot pass the verification are not used and ranked. The server $S$ then selects the top $N$ clients based on their ranked performance metrics into the candidate set $\mathcal{C}$ and notifies the respective clients in this set to upload their updated local models.
@@
    The server $S$ collects the updated local models sent from the $N$ clients in the candidate set $\mathcal{C}$. The server first computes the hash of the updated local model sent by each selected client in step 5: $\widetilde{h}_{k}^{t}=\text{hash}(w_{k}^{t})$, and compares the computed hash with the hash, $h_{k}^{t}$, sent by the client in step 3. If the two hashes match, the server $S$ will labels this client as a valid candidate; if the two hashes differ, the server $S$ will remove this client $C_{k}$ from the candidate set $\mathcal{C}$. After this hash verification operation for each client $C_{k}\in\mathcal{C}$, the server $S$ aggregates these updated local models of the clients still in the candidate set $\mathcal{C}$ according to the predetermined rule to derive the updated global model $w^{t}$ for the $t$-th round and then distribute $w^{t}$ to all the clients.
The above processes of Veri-CS-FL iterate continuously until the final global model converges or the predetermined number of epochs is reached. Some remarks about Veri-CS-FL are given below.
0: Each client $C_{k}$ possesses the proving key $\mathsf{PK}$ of zk-SNARK and the local dataset $D_{k}$ , $\forall k\in\{1,2,\cdots,K\}$ while the server $S$ holds the verifying key $\mathsf{VK}$ of zk-SNARK and the small and clean dataset $D_{s}$.
@@
9:   generates $\pi_{k}^{t}$ of $w_{k}^{t}$: $\pi_{k}^{t}\leftarrow\mathsf{PROVE}(\mathsf{PK},x,W)$ with $h_{k}^{t}=\text{hash}(w_{k}^{t})$, $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$ and $W=w_{k}^{t}$;
10:   transmits $\pi_{k}^{t}$ and $h_{k}^{t}$ to the server $S$.
@@
14:     verifies all the proofs: $b\leftarrow\mathsf{VERIFY}(\mathsf{VK},x,\pi)$ with $x=\left\{{w_{s}^{t},h_{k}^{t}}\right\}$ and $\pi=\pi_{k}$;
@@
22:     compares $\widetilde{h}_{k}^{t}$ with $h_{k}^{t}$ (line 9): if $\widetilde{h}_{k}^{t}\neq h_{k}^{t}$, removes the corresponding $C_{k}$ from $\mathcal{C}$, $\forall k\in\{1,2,\cdots,N\}$.
@@
    In Veri-CS-FL, each client needs to submit the performance metric of their local model to the server. The server then uses the performance metrics to select local models with satisfied performances, and uses the training sample sizes as the weights in the weighted aggregation method of global models. There are many different performance metrics that can be used to characterize the performance of a local model. For example, the loss function of the model on the test dataset, the KL divergence of the output predicted by the model on the test dataset. In Veri-CS-FL, we propose to adopt the cosine similarity between the benchmark model and the local model [7] as the performance metric: $\text{cos}_{k}^{t}\triangleq{\text{cos}}\left(w_{k}^{t},w_{s}^{t}\right)=\frac{w_{k}^{t}\cdot w_{s}^{t}}{|w_{k}^{t}||w_{s}^{t}|}$. When utilizing the benchmark model as a reference, the larger the cosine similarity between the local model and the benchmark model, the better the performance of the local model. The computation of cosine similarity can also be easily translated to the arithmetic circuit of zk-SNARK, enabling the computation of the zero-knowledge proof for it.
@@
    By generating the zero-knowledge proofs for the performance metrics of local models at the client-sides, and validating the proofs at the server side, Veri-CS-FL can achieve the following advantages: i) Communication-efficient: the local models with poor performances will not be communicated between the clients and the server. This can save much uplink communication bandwidth, particularly when the models are large. ii) Computation-efficient: the server has no need to re-run all the local models to evaluate their performances. The verification of zero-knowledge proofs in zk-SNARK is very efficient, e.g., the zk-SNARK algorithm of Groth16 [4] has a fixed verification complexity and fixed proof size regardless of the sizes of public inputs and private witness. iii) Trusted collaboration: The client transmits the performance metrics and its zero-knowledge proofs to the server, and if the server feels that its updated local model has satisfactory performance, it asks the client for the updated local model. As a result, the client and server can agree on the model training contribution. It prevents cases where after the client passes the updated local model to the server, the server refuses to admit that the client provided a local model with satisfactory performance in this round of training, and maliciously devalues the actual contribution of the client.
@@
IV-C Security and Privacy Analysis
In this subsection, we analyze the security properties of Veri-CS-FL, focusing on completeness, soundness, zero-knowledge, and resistance to poisoning attacks.
IV-C1 Completeness
For any honest client $C_{k}$ that correctly executes the local training and metric calculation protocol, the generated proof $\pi_{k}$ will always pass the verification $\mathsf{VERIFY}(\mathsf{VK},x,\pi_{k})=1$. This ensures that legitimate contributions are always recognized by the server.
IV-C2 Soundness
The soundness property ensures that a malicious client cannot generate a valid proof for a false statement. Specifically, if a client submits a high performance metric but has not actually trained a model that yields such a metric (or has not trained a model at all), it is computationally infeasible to generate a proof $\pi$ that satisfies the verification algorithm. This prevents clients from free-riding or falsifying their contributions.
IV-C3 Zero-Knowledge
The zero-knowledge property guarantees that the proof $\pi_{k}$ reveals no information about the witness $W$ (the local model weights $w_{k}^{t}$) other than the validity of the statement (the cosine similarity metric). The server learns only the metric and the hash of the model, ensuring that the model parameters remain private until the client is selected for aggregation. Notably, the only information exposed is a single scalar value (the cosine similarity score), which is insufficient to reconstruct the high-dimensional model parameters or infer the underlying training data distribution. The zero-knowledge property preserves the privacy of FL.
IV-C4 Resistance to Poisoning Attacks
Veri-CS-FL inherently resists model poisoning attacks through its verifiable client selection mechanism. By filtering models based on their cosine similarity to a benchmark model trained on clean data, the system effectively excludes malicious updates that deviate significantly from the benign optimization direction. Unlike standard FL, where the server aggregates all models (including potentially poisoned ones), Veri-CS-FL verifies the quality before aggregation.
IV-C5 Impact on Heterogeneous Data
A potential concern is whether filtering based on a benchmark model restricts the learning of unique features from outlier datasets. Veri-CS-FL mitigates this by using cosine similarity, which measures the alignment of the update direction rather than the specific parameter values. Valid updates from non-IID clients, even if they contain unique features, generally align with the global optimization direction (positive similarity). In contrast, malicious or purely noisy updates typically exhibit orthogonality or opposition. Furthermore, by selecting the top $N$ contributors rather than applying a rigid threshold, the system adapts to the available pool of updates, preserving the most constructive contributions even in heterogeneous environments.